# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v1

@@ III Results and Discussion
In this section, we use the results of DNS to provide insight into the effect of buoyancy on the flow field behind the cylinder. Based on the instantaneous and time-averaged results, we discuss the physical mechanisms for the suppression of vortex shedding, vorticity inversion, and the far-field unsteadiness. The Stuart-Landau model has been employed to accurately estimate the critical Richardson number for the suppression of vortex shedding in the near-field. We have also discussed the effect of buoyancy on the integral parameters of the flow field. The results obtained are only applicable to the streamwise length of the domain used in the current study. The choice of domain length influences the observed transitions, and it is possible that a longer domain could reveal additional flow physics. To achieve statistical independence, we set the total averaging time (T) to different values for different calculations. It is found that the results obtained do not vary more than $2\%$ between two different averaging times. Therefore, it is expected that the presented results are statistically independent.
@@ III.1 Instantaneous flow
Instantaneous vorticity and temperature contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$ are presented in Figures 3 and 5 for the full domain and Figures 4 and 6 in the vicinity of the cylinder, respectively. All instantaneous contours are phase-synchronized to ensure uniformity in flow visualization. The instantaneous vorticity contours (see Figure 3) reveal the presence of vortex shedding and the Bénard-von Kármán vortex street in the near-field for Ri $=0.0$ (forced convection). With the absence of the buoyancy-driven flow, the flow field resembles that of forced flows, showing negative vorticity in the left half and positive vorticity in the right half, which spans the entire domain downstream. The two oppositely-signed shear layers attached to the cylinder cross the centerline alternatively during a shedding cycle and interact with each other, leading to the shedding of vortices. In the far-field, the flow is steady at Ri $=0.0$. When Ri is increased to 0.2, it is seen that the sign of vorticity starts out as that of a forced flow but slowly inverts beyond a critical distance downstream and persists with this inverted sign for the rest of the domain beyond the critical distance. Additionally, the far-field flow is no longer steady as it was in the Ri $=0.0$ case. At higher Ri (0.4-0.7), the inversion of vorticity takes place closer to the cylinder, and the width between the two shear layers increases downstream. The shape of the shed vortices also differs greatly with increasing Ri, which requires further investigation. For Ri $\geq$ 0.7, the near-field flow becomes steady with the suppression of vortex shedding by buoyancy, and a large-scale undulation of the shear layers is seen in the far-field region. The shear layer becomes discontinuous, leading to the formation of isolated patches of vorticity due to the instability developing within the stronger shear layer. As Ri increases, the inception point of the far-field unsteadiness moves closer to the cylinder. The instantaneous vorticity in the cylinder vicinity (see Figure 4) shows regular vortex shedding at Ri $=0.0$ with vortices of opposite sign being shed alternatively into the wake. With increasing Ri, the transverse extent of the interaction of the positively-signed shear layer with the oppositely-signed shear layer reduces gradually, appearing as though their mutual interaction is being inhibited. For Ri $>$ 0.7, there is no interaction between the oppositely-signed shear layers, and they form a separation bubble, similar to forced flows below the critical Re at which the transition to unsteadiness occurs. Vortex shedding does not occur for Ri $>$ 0.7 and the point of inversion of vorticity comes so close to the cylinder that it can be visualized in the near-field plots of vorticity as well (see Figures 4(e) and 4(f)). Isotherms reveal the presence of vortex shedding in the near-field for Ri $=0.0$ and the far-field is steady, as confirmed through the examination of signals of temperature, even though small oscillations are noticeable in the isotherm (see Figure 5). At Ri $=0.2$, the undulation of shear layers extends to Y $=65$ and the complete flow field is unsteady up to Ri $=0.6$. The shape of the isotherm structures continues to change when Ri is increased, due to the increased strength of the buoyancy force and the constant strength of the inertial force. With further increase in Ri, the temperature in the near-field is steady, and the far-field reveals the presence of structures similar to those seen in plumes, as shown by Kimura and Bejan. [31] The near-field plots of temperature show similar behavior as the vorticity contours (see Figure 6). The undulation of the isotherms is significantly lower at higher Ri, and for Ri $\geq$ 0.7, the structures reveal a steady flow in the near-field.
@@
Figure 3: Instantaneous vorticity contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
@@
Figure 4: Instantaneous vorticity contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$; ($\overline{\omega}_{Z,min},\overline{\omega}_{Z,max},\Delta\overline{\omega}_{Z})\equiv(-5,5,0.4)$.
@@
Figure 6: Instantaneous temperature contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$; ($\overline{\theta}_{min},\overline{\theta}_{max},\Delta\overline{\theta})\equiv(0.01,1,0.041)$.
@@
Figure 7: Streamwise (a, b) and transverse (c, d) velocity signals in the near- and far-field for Ri $=0.6$ and Ri $=0.7$, along with (e) drag and lift coefficient signals. The near-field signals reveal the suppression of vortex shedding and the far-field signals show the far-field unsteadiness, while the force coefficient signals characterize the near-field flow excellently.
To confirm the behavior shown by the flow field in the contours, we have plotted time signals of the streamwise and transverse velocity in the near-field (Y $=11.5$) and far-field (Y $=73.5$), which are shown in Figure 7 for Ri $=0.6$ and Ri $=0.7$. For Ri $=0.6$, streamwise and transverse velocity plots in the near-field reveal unsteady, periodic behavior, and for Ri $=0.7$, steady behavior without any periodicity is observed (see Figures 7(a) and 7(c)). Far-field streamwise and transverse velocity plots show the opposite behavior, with mild unsteadiness in the case of Ri $=0.6$ associated with increasing plume-like behavior and large-scale unsteadiness for Ri $=0.7$ associated with the onset of the far-field plume-like unsteadiness (see Figures 7(b) and 7(d)). There is a large difference in the amplitude of the signals between Ri $=0.6$ and Ri $=0.7$, which confirms that the unsteadiness in the far-field is not due to the shedding of vortices in the near-field, the strength of which dissipates significantly due to viscous diffusion as it reaches the far-field, which can be seen in the Ri $=0.6$ signals. In both regions (near-field and far-field), the streamwise velocity signal has a frequency that is twice that of the transverse velocity signal. Interestingly, in the near-field, U exhibits multiple frequencies, while V has a unique frequency. In the far-field, V exhibits multiple frequencies, while U has a unique frequency, demonstrating an opposite behavior to that of the near-field. It appears that such behavior is characteristic of the wake nature of the near-field and the plume nature of the far-field; however, this requires further investigation. The plots of lift and drag coefficient (see Figure 7(e)) show similarities with the near-field velocity signals, suggesting that the near-field flow is unsteady for Ri $=0.6$ and steady for Ri $=0.7$, supporting the suppression of vortex shedding in the near-field. This makes the force coefficients excellent predictors of the nature of the near-field flow. The skin friction drag and Nusselt number are analogous quantities in that they are both proportional to the gradient of velocity and temperature, respectively. Similar to the variation in friction drag, which is negligible, the fluctuation in Nu is also very small, even though its time-averaged value is very large. Therefore, Nu has not been plotted as a function of time, as there are negligible fluctuations in Nu for all Ri.
Table 7: Variation of Strouhal number (St) with Richardson number (Ri) in the near-field (Y $=11.5$) and far-field (Y $=73.5$). The transverse velocity signal along the centerline (X $=0$) has been used for the calculation of Strouhal number.
Richardson number (Ri)
Strouhal number
@@
Table 7 shows the variation of Strouhal number (St) with Ri. The transverse velocity signal along the centerline (X $=0$) is used to compute the Strouhal number at two different streamwise locations, Y $=11.5$ and Y $=73.5$ which fall in the near-field and far-field, respectively. St shows negligible variation between the near-field and far-field for Ri $<$ 0.7. Above the critical Ri, the near-field flow has no periodicity and the corresponding St is zero. The flow in the wake is accelerated due to the buoyancy force, which is more significant at elevated Richardson numbers, resulting in increased Strouhal number with increasing Ri.
III.2 Stuart-Landau Analysis
In DNS, we are able to identify the critical Ri correct to a certain precision due to the discrete increments of the control parameter (here, Ri) taken during calculations. The Stuart-Landau model enables accurate identification of the critical Ri. This model is well-known for describing the behavior of nonlinear dynamical systems near the Hopf bifurcation. The derivation of this equation from the hydrodynamic equations has been provided in many previous works. [34, 35, 36, 37] The Landau model has been shown to accurately describe the transition to periodicity in the wake of flows past cylinders by several authors in the literature. [3, 38, 39] It has been shown earlier by Dushe [17] that the two transitions obtained for mixed convection past a square cylinder at $\alpha=0^{\circ}$ are both Hopf bifurcations. Dushe also showed that the Stuart-Landau model is suitable for the prediction of the critical Ri for vortex shedding suppression in the wake of mixed convective flows past cylinders. The Stuart-Landau equation describes the perturbation near the bifurcation as follows:
[…]
Where $A$ is any characteristic complex amplitude associated with the fundamental frequency, $\sigma$ is the linear growth rate, and $l$ is the first Landau constant. It can be shown that the equilibrium amplitude ($A_{e}$) of oscillations near the bifurcation varies with the control parameter (Ri) as:
$A_{e}^{2}=C(Ri-Ri_{c})$ (11)
Where $Ri_{c}$ is the critical Richardson number and C is a constant. Equilibrium amplitudes of the characteristic quantity around the (approximate) critical Richardson number can be plotted and fitted to this equation. The intercept of the best-fit line with the Ri axis yields the critical Richardson number.
For selecting the most appropriate characteristic quantity, we have plotted the mean-to-peak amplitudes of different quantities ($C_{L}$, $C_{D}$, $v$, $u$, $\theta$) against Ri, below the approximate critical Ri and fitted a linear curve to the data. Based on this, we have found that the lift coefficient produces the best fit in terms of the coefficient of determination ($R^{2}$) as well as closeness to the critical Ri predicted by DNS. Therefore, we have used the lift coefficient signal for estimation of the near-field transition point. The Landau model is applicable only for those Ri very close to the critical Ri as it is a linearized model. Therefore, we have considered the equilibrium amplitudes of the lift coefficient for Ri in the range (0.64-0.68) in increments of 0.01. The square of the equilibrium amplitude ($A_{e}^{2}$) of the lift coefficient signal has been plotted against Ri along with the best-fit line in Figure 8. The critical Ri determined from the Landau model comes out to be 0.681. The corresponding value obtained by Arif and Hasan [28] is 0.78 which is different from our result. This may be due to a different discretization in the computational method or the short domain taken by the authors.
Figure 8: Variation of $A_{e}^{2}$ of the lift coefficient signal with Richardson number.
III.3 Time-averaged flow
@@
Time-averaged vorticity and temperature contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$ are presented in Figures 9 and 11 in the full domain and Figures 10 and 12 in the vicinity of the cylinder, respectively. The time-averaged vorticity contours in the complete domain (see Figure 9) reveal that at Ri $=0.0$, the strength of the vorticity is high in the near-field but diminishes in the far-field, while for Ri $>$ 0.0, the strength of the vorticity is weak in the near-field but grows on moving downstream, showing the opposite nature of the forced convective vorticity and the natural convective vorticity. The forced convective vorticity is introduced due to the shear layer separation at the lateral corners of the cylinder, while the natural convective vorticity is introduced due to the buoyancy-driven flow, which causes a jetting effect. The interplay of these opposing vorticities is the cause of many physical phenomena that we have observed in this study and will be discussed in detail later. Beyond Ri $=0.7$, the strength of the vorticity is significantly higher than that of the cases below Ri $=0.7$, and the oppositely-signed vortices remain attached up to the far-field, which was not the case at lower Ri. Further, the two vortices separate suddenly in the far-field at a point (around Y $=65$) which also coincides with the point of onset of the far-field unsteadiness. All positive Ri cases display vorticity inversion as seen in the instantaneous contours earlier. In the near-field plot of vorticity, the width between the shear layers is seen to reduce with increasing Ri, similar to the instantaneous case. Beyond Ri $=0.7$, the inversion can be seen in the near-field, and the vortices form a steady recirculation bubble, whose length decreases with increasing Ri, and these cases resemble the instantaneous plots as the flow is steady in the near-field. Time-averaged temperature contours (see Figure 11) show a widening of the plume in the near and intermediate fields for Ri $<$ 0.7. Beyond 0.7, the plume is narrow in the near-field, and there is a sharp increase in the plume width in the far-field, associated with the onset of the far-field unsteadiness. Such behavior is also observed in the near-field temperature contours (see Figure 12), where the plume is wide in the near-field for low Ri, and beyond Ri $=0.7$, the plume is narrow.
@@
Figure 10: Time-averaged vorticity contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$; ($\overline{\omega}_{Z,min},\overline{\omega}_{Z,max},\Delta\overline{\omega}_{Z})\equiv(-4,4,0.53)$.
@@
Figure 12: Time-averaged temperature contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$, ($\overline{\theta}_{min},\overline{\theta}_{max},\Delta\overline{\theta})\equiv(0.01,1,0.041)$
@@
Figure 13 shows the centerline streamwise velocity and centerline temperature along the streamwise length of the domain for different Ri. Figure 13(a) reveals the buoyancy-driven jetting effect, which becomes more pronounced as one moves further downstream. There is a deficit of momentum in the immediate wake of the cylinder due to shear layer separation at the lateral corners. However, the presence of the buoyancy-driven natural convective flow counteracts this deficit and accelerates the flow behind the cylinder, leading to a recovery and, eventually, an excess of momentum beyond the cylinder, which causes the flow field to show jet-like characteristics far downstream, in the time-averaged sense, while it exhibits a wake-like nature in the near-field. The average streamwise velocity ($V_{avg}$) at any particular downstream location Y is equal to the free-stream velocity ($V_{\infty}$ = 1, in the present case). Therefore, if the centerline velocity is lower than $V_{avg}$, there is a momentum deficit, and if the centerline velocity exceeds $V_{avg}$, there is a momentum excess at that location. For the purpose of visualizing the location of momentum recovery and subsequent excess, $\overline{V}$ = 1 has been plotted in Figure 13(a). The centerline velocity variation reveals that the momentum deficit persists in the forced convection case throughout the domain, despite the recovery of some momentum later downstream. For positive Ri, the momentum recovery is aided by the buoyancy-driven jet, and the recovery distance decreases with increasing Ri, occurring at Y $=20$ for Ri $=0.2$ and around Y $=5$ for Ri $\geq$ 0.6. Further downstream, for Ri $\geq$ 0.7, there is a decrease in velocity in the far-field, which occurs around the same distance as the onset of the far-field unsteadiness. The inset in the same graph shows the centerline velocity in the immediate wake of the cylinder (0.5 $\leq$ Y $\leq$ 1.75) which provides insight on the recirculation length. Starting from the base of the cylinder (Y $=0.5$), there is a negative streamwise velocity due to the recirculation of flow as it turns towards the cylinder (in the time-averaged sense). However, the velocity quickly becomes positive after the recirculation region. The recirculation length shows a nonlinear variation, with Ri $=1$ and Ri $=0.2$ having the shortest recirculation length, followed by Ri $=0.0$, 0.6 and 0.7, respectively. Figure 13(b) shows the variation of centerline temperature with streamwise distance, revealing a dissipating trend of temperature with downstream distance. The centerline temperature is highest for higher Ri close to the cylinder; however, in the far-field, the behavior is the opposite, and higher Ri have the lowest temperatures, albeit the difference is marginal in the far-field.
@@
Nusselt number is a function of the spatial coordinates as well as time. Therefore, it may be averaged in space, time, or both. $\overline{Nu}_{local}$ refers to the time-averaged local Nusselt number at a particular location on the cylinder, while $\overline{Nu}_{surface}$ refers to the temporally and spatially averaged Nusselt number for a particular surface on the cylinder, and $\overline{Nu}$ refers to the temporally and spatially averaged Nusselt number for the entire cylinder. Figure 14 shows the variation of time-averaged Nusselt number for different cylinder surfaces and local Nusselt number along the cylinder surface for various Ri, along with the nomenclature of the cylinder surfaces (AB-DA). For flow-facing surfaces (AB and DA), the mean Nusselt number increases with increasing Ri, whereas for wake-facing surfaces (BC and CD), the mean Nusselt number decreases with increasing Ri. There is a sharp decrease in the Nusselt number at Ri $=0.7$ for the wake-facing surfaces, and subsequently, in the Nusselt number for the full cylinder. The shedding phenomenon helps heat transfer by increasing mixing in the fluid by the mutual interaction of the two shear layers. When vortex shedding is suppressed, the momentum transport is inhibited as the mutual interaction of the shear layers is no longer active. Therefore, the Nusselt number sharply decreases for the wake-facing surfaces at Ri $=0.7$. The Nusselt number for the surfaces AB and DA, as well as that for surfaces BC and CD, are equal due to the symmetric nature of the time-averaged flow about the centerline. The local Nusselt number is the highest for the vertex A, which directly faces the forced flow, and it is considerably lower for the wake-facing surfaces compared to the flow-facing surfaces, due to the recirculation region having very weak streamwise velocities, causing decreased heat transfer from the wake-facing surfaces. The Nusselt number reaches a local maxima at the lateral corners of the cylinder (B and D), due to the acceleration of flow at these corners.

### v3

@@ III Results and Discussion
In this section, we have used the results of DNS to provide insight into the effect of buoyancy on the flow field behind the cylinder. Based on the instantaneous and time-averaged results, we have discussed the physical mechanisms for the suppression of vortex shedding, vorticity inversion, and far-field unsteadiness. We have also discussed the effect of buoyancy on the integral parameters of the flow field. It is to be noted that the results presented in this study are only applicable to the streamwise length of the domain used in the current study. The choice of domain length influences the observed transitions, and a longer domain could reveal additional flow physics. We set the total averaging time (T) to two different values for different calculations to achieve statistical independence. The values taken for T are 200 and 300, which correspond to 36 and 54 shedding cycles at Ri = 0.0, respectively (which has the highest shedding time period due to having the lowest Strouhal number as seen in Table 6). It is found that the results obtained do not vary more than $1\%$ between two different averaging times. Therefore, it is expected that the presented results are statistically independent.
@@ III.1 Instantaneous Flow
Instantaneous vorticity contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$ in the full domain, near-field, and far-field are presented in Figures 3, 4 and 5 respectively. Instantaneous temperature contours in the full domain and near-field are shown in Figures 6 and 7 respectively. All instantaneous contours are phase-synchronized (phase in the shedding cycle when the lift coefficient is minimum) to ensure uniformity in flow visualization.
@@
The instantaneous vorticity contours (see Figure 3) reveal the presence of the Bénard-von Kármán vortex street, signifying the shedding of vortices in the near-field for Ri $=0.0$ (forced convection). There is negative vorticity in the left and positive vorticity in the right half, which spans the entire domain downstream. The two oppositely-signed shear layers attached to the cylinder interact, leading to the shedding of vortices. The flow becomes steady at the far-field due to viscous diffusion. When Ri is increased to a non-zero positive value, it is seen that the sign of vorticity starts as that of a forced flow but slowly inverts beyond a critical distance downstream and persists with this inverted sign for the rest of the domain beyond the critical distance. Additionally, the far-field flow is no longer steady, which was the case for Ri $=0.0$. At higher Ri (0.4-0.7), the inversion of vorticity takes place closer to the cylinder, and the width between the two shear layers increases downstream. The shape of the shed vortices also differs greatly with increasing Ri, which requires further investigation. For Ri $\geq$ 0.7, the near-field flow becomes steady with the suppression of vortex shedding due to the effect of buoyancy, and a large-scale undulation of the shear layers is seen in the far-field region. The shear layer becomes discontinuous, forming isolated patches of vorticity due to the instability developing within the stronger shear layer. As Ri increases, the inception point of the far-field unsteadiness moves closer to the cylinder. The instantaneous vorticity in the cylinder vicinity (see Figure 4) shows regular vortex shedding at Ri $=0.0$ with vortices of the opposite sign being shed alternatively into the wake. With increasing Ri, the transverse extent of the interaction of the positively-signed shear layer with the oppositely-signed shear layer reduces gradually, appearing as though their mutual interaction is being inhibited. For Ri $>$ 0.7, there is minimal interaction between the oppositely-signed shear layers, and they form a separation bubble, similar to forced flows of a sub-critical Re at which the transition to unsteadiness occurs. Vortex shedding from the cylinder does not happen for Ri $>$ 0.7, and the point of inversion of vorticity comes so close to the cylinder that it can be visualized in the near-field plots of vorticity as well (see Figures 4(e) and 4(f)).
@@
Figure 3: Instantaneous vorticity contours in the complete domain for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
@@
Figure 4: Instantaneous vorticity contours in the near-field for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$; ($\omega_{baro,min},\omega_{baro,max},\Delta\omega_{baro})\equiv(-0.1,0.1,0.01)$. The flood represents total vorticity, while the lines represent baroclinic vorticity.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
Figure 5: Instantaneous vorticity contours in the far-field for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$; ($\omega_{baro,min},\omega_{baro,max},\Delta\omega_{baro})\equiv(-0.05,0.05,0.005)$. The flood represents total vorticity while the lines represent baroclinic vorticity.
The inversion of vorticity has been reported in the literature for isothermal confined flows [14, 15, 16, 17], where the interplay between several sources of vorticity in the flow causes the reversal of vorticity in the wake of the cylinder. In unconfined flows, a reversal of the Bénard-von Kármán vortex street takes place when the obstacle (bluff body or aerofoil) is imparted a periodic motion. Under appropriate conditions of the shape of the undulating body or the control parameter, the shed vortices cross the centerline and switch their position after travelling downstream, which manifests as a jet profile with a momentum excess in the time-averaged velocity plot, compared to the original wake profile with a momentum deficit [34, 35, 36]. The vorticity inversion seen in our case resembles such structures. However, the shed vortices do not physically convect to the opposite side in this case. Rather, there is an inversion of vorticity due to the interplay between the vorticity fields due to forced convection and natural convection, which will be discussed in a later section.
The forced convective component of the flow causes shear layer separation at the transverse corners of the cylinder, causing a deficit of momentum in the wake. On the other hand, the natural convective component causes jet-like behavior downstream, competing with the momentum deficit caused by the wake, with the respective vorticities signed oppositely to each other. The opposing effects of these two flow components lead to a vorticity inversion. This inversion does not occur for Ri $=0.0$ due to the absence of buoyancy at this Ri. For positive Ri, vorticity inversion occurs downstream at an inception point, which comes further upstream as Ri is increased (see Figure 15) due to the increased buoyancy (see Figures 3 and 9).
Isotherms also reveal the presence of vortex shedding in the near-field for Ri $=0.0$ and the steady far-field, as confirmed by examining temperature signals. At Ri $=0.2$, the undulation of shear layers extends to Y $=65$, and the complete flow field is unsteady up to Ri $=0.6$. The shape of the isotherm structures continues to change when Ri is increased due to the increased strength of the buoyancy force and the constant strength of the inertial force. With a further increase in Ri, the temperature in the near-field becomes steady, while the far-field reveals the presence of structures similar to those seen in plumes, as shown by Kimura and Bejan [31]. The near-field plots of temperature show similar behavior as the vorticity contours (see Figure 7). The undulation of the isotherms becomes significantly lower with increasing Ri, and for Ri $\geq$ 0.7, the structures reveal a steady flow and thermal field in the near-field.
@@
Figure 7: Instantaneous temperature contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
@@
Figure 8: Streamwise (a, b) and transverse (c, d) velocity signals in the near- and far-field for Ri $=0.6$ and Ri $=0.7$, along with (e) drag and lift coefficient signals. The near-field signals reveal the suppression of vortex shedding, and the far-field signals show the far-field unsteadiness. In contrast, the force coefficient signals characterize the near-field flow excellently.
To confirm the behavior shown by the flow field in the contours, we have plotted time signals of the streamwise and transverse velocity in the near-field (Y $=11.5$) and far-field (Y $=73.5$), which are shown in Figure 8 for Ri $=0.6$ and Ri $=0.7$. For Ri $=0.6$, streamwise and transverse velocity plots in the near-field reveal unsteady, periodic behavior, and for Ri $=0.7$, a steady behavior without any periodicity is observed (see Figures 8(a) and 8(c)). Far-field streamwise and transverse velocity plots show the opposite behavior, with negligible unsteadiness in the case of Ri $=0.6$ but a large-scale unsteadiness for Ri $=0.7$ associated with the onset of the far-field plume-like unsteadiness (see Figures 8(b) and 8(d)).
The plots of lift and drag coefficient (see Figure 8(e)) show similarities with the near-field velocity signals, suggesting that the near-field flow is unsteady for Ri $=0.6$ and steady for Ri $=0.7$, supporting the suppression of vortex shedding in the near-field. This makes the force coefficients excellent predictors of the nature of the near-field flow.
Table 6: Variation of Strouhal number (St) with Richardson number (Ri) in the near-field (Y $=11.5$) and far-field (Y $=73.5$). The transverse velocity signal along the centerline (X $=0$) has been used to calculate the Strouhal number.
Richardson number (Ri) Strouhal number (St)
@@
Table 6 shows the variation of Strouhal number (St) with Ri. The transverse velocity signal along the centerline (X $=0$) is used to compute the Strouhal number at two different streamwise locations, Y $=11.5$ and Y $=73.5$, which fall in the near-field and far-field, respectively. St shows negligible variation between the near-field and far-field for Ri $<$ 0.7. Above the critical Ri, the near-field flow has no periodicity, and the corresponding St is zero. The flow in the wake is accelerated due to the buoyancy force, which is more significant at elevated Richardson numbers, resulting in increased Strouhal number with increasing Ri.
III.2 Time-Averaged Flow
@@
Time-averaged vorticity and temperature contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$ are presented in Figures 9 and 11 in the full domain and Figures 10 and 12 in the vicinity of the cylinder, respectively. The time-averaged vorticity contours in the complete domain (see Figure 9) reveal that at Ri $=0.0$, the strength of the vorticity is high in the near-field but diminishes in the far-field. In contrast, for Ri $>$ 0.0, the strength of the vorticity is weak in the near-field but grows on moving downstream, showing the opposite nature of the forced convective vorticity and the natural convective vorticity. The forced convective vorticity is introduced due to the shear layer separation at the transverse/lateral corners of the cylinder, while the natural convective vorticity is introduced due to buoyancy, which causes a jet-like effect in the wake. The interplay of these opposing vorticities is the cause of many physical phenomena that we have observed in this study and will be discussed in detail later. Beyond Ri $=0.7$, the strength of the vorticity is significantly higher than that of the cases below Ri $=0.7$, and the oppositely-signed vortices remain attached up to the far-field, which was not the case at lower Ri. Further, the two vortices are found to separate suddenly in the far-field at a point (around Y $=65$) which also coincides with the point of onset of the far-field unsteadiness. All positive Ri cases display vorticity inversion, as seen in the instantaneous contours earlier. In the near-field plot of vorticity, the width between the shear layers is seen to reduce with increasing Ri, similar to the instantaneous case. Beyond Ri $=0.7$, the inversion can be seen in the near-field, and the vortices form a steady recirculation bubble, whose length decreases with increasing Ri. These cases resemble the instantaneous plots as the flow is steady in the near-field. Time-averaged temperature contours (see Figure 11) show a widening of the plume in the near and intermediate fields for Ri $<$ 0.7. Beyond Ri=0.7, the plume is narrow in the near-field, and there is a sharp increase in the plume width in the far-field, associated with the onset of the far-field unsteadiness. Such behavior is also observed in the near-field temperature contours (see Figure 12), where the plume is wide in the near-field for low Ri, and beyond Ri $=0.7$, the plume is narrow.
@@
Figure 10: Time-averaged vorticity contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
@@
Figure 12: Time-averaged temperature contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
@@
Figure 13 shows the centerline streamwise velocity and centerline temperature along the streamwise length of the domain for different Ri. Figure 13(a) reveals the buoyancy-driven jetting effect, which becomes more pronounced as one moves further downstream. The average streamwise velocity ($V_{avg}$) at any particular downstream location Y is equal to the free-stream velocity ($V_{\infty}$ = 1.0, in the present case). It is to be noted that the wake has a momentum deficit when the centerline streamwise velocity is lower than $V_{avg}$ and becomes a wake with an excess momentum if the centerline velocity exceeds $V_{avg}$ at that location. There is a deficit of momentum in the immediate wake of the cylinder due to shear layer separation at the front lateral corners. However, the presence of the natural convective flow counteracts this deficit and accelerates the flow behind the cylinder, leading to a faster recovery and, eventually, an excess of momentum beyond the cylinder exhibiting jet-like characteristics far downstream, while the flow immediately downstream of the cylinder shows a wake-like behavior in the near-field. For the purpose of visualizing the location of momentum recovery and subsequent excess, $\overline{V}$ = 1.0 has been plotted in Figure 13(a). The centerline velocity variation reveals that the momentum deficit persists in the forced convection case throughout the domain despite the recovery of some momentum later downstream. For positive Ri, the momentum recovery is aided by the buoyancy-driven jet, and the recovery distance decreases with increasing Ri, occurring at Y $=20$ for Ri $=0.2$ and around Y $=5$ for Ri $\geq$ 0.6. Further downstream, for Ri $\geq$ 0.7, there is a decrease in velocity in the far-field, which occurs around the same distance as the onset of the far-field unsteadiness. This phenomenon is attributed to the widening of the wake caused by the unsteadiness of the flow beyond this location. The inset in the same graph shows the centerline velocity in the immediate wake of the cylinder (0.5 $\leq$ Y $\leq$ 1.75), providing an insight into the recirculation length. Starting from the base of the cylinder (Y $=0.5$), there is a negative streamwise velocity due to the recirculation of flow as it turns towards the cylinder (in the time-averaged sense). However, the velocity quickly becomes positive after the recirculation region. The recirculation length shows a nonlinear variation, with Ri $=1$ and Ri $=0.2$ having the shortest recirculation length, followed by Ri $=0.0$, 0.6 and 0.7, respectively. Figure 13(b) shows the centreline variation of temperature with streamwise distance, revealing a reducing trend of temperature with downstream distance. The centerline temperature is highest for higher Ri close to the cylinder; however, in the far-field, the behavior is the opposite, and higher Ri have lower temperatures, albeit the difference is marginal in the far-field.
@@
Nusselt number is a function of the spatial coordinates as well as time. Therefore, one can average the Nusselt number in space, time, or both. $\overline{Nu}_{local}$ refers to the time-averaged local Nusselt number at a particular location on the cylinder, while $\overline{Nu}_{surface}$ refers to the temporally and spatially averaged Nusselt number for a particular surface on the cylinder, and $\overline{Nu}$ refers to the temporally and spatially averaged Nusselt number for the entire cylinder. Figure 14 shows the variation of time-averaged Nusselt number for different cylinder surfaces and local Nusselt number along the cylinder surface for various Ri, along with the nomenclature of the cylinder surfaces (AB, BC, CD, and DA). For flow-facing surfaces (AB and DA), the mean Nusselt number increases with increasing Ri, whereas for wake-facing surfaces (BC and CD), the mean Nusselt number decreases with increasing Ri. There is a sharp decrease in the Nusselt number at Ri $=0.7$ for the wake-facing surfaces and, subsequently, in the Nusselt number for the full cylinder. The shedding phenomenon helps in higher heat transfer due to increasing momentum exchange between the thermal boundary layers at various walls and the separating shear layers. When vortex shedding is suppressed, the momentum transport is inhibited as the mutual interaction is no longer active. Therefore, the Nusselt number sharply decreases for the wake-facing surfaces at Ri $=0.7$. The Nusselt number for the surfaces AB and DA, as well as that for surfaces BC and CD, are equal due to the symmetric nature of the time-averaged flow about the centerline. The local Nusselt number is the highest for the vertex A, which directly faces the forced flow, and it is considerably lower for the wake-facing surfaces compared to the flow-facing surfaces due to the recirculation region having very weak streamwise velocities, causing decreased heat transfer from the wake-facing surfaces. The Nusselt number reaches a local maxima at the lateral corners of the cylinder (B and D), due to the acceleration of flow at these corners.